OpenStudy (anonymous):

Any help is greatly appreciated please and thank- you Part 1: Create your own quadratic equation that cannot be solved by factoring, but can be solved using the quadratic formula. Identify the values of a, b, and c, and find the solutions using the quadratic formula. Show all work to receive credit. Part 2: Using complete sentences, explain how you know that the equation from Part 1 cannot be solved by factoring, but can be solved by using the quadratic formula.

OpenStudy (anonymous):

Part 1: 2x^2 - 3x - 4 = 0 b^2 - 4ac = (-3)^2 - 4(2)(-4) = 9 + 32 = 41 a = 2, b = -3, c = -4 From the Quadratic Formula, x ≈ 2.35 or x ≈ -0.85 Part 2: Because 41 is not a perfect square, we know the equation is not factorable.

OpenStudy (anonymous):

thanks but can you show me how to use the quadratic formula for this question

OpenStudy (anonymous):

Sorry, I leave using the Quadratic Formula for you to do. It is hard to do on the computer, but easy to do by hand.